Question 1.

Dandy’s manufacturers Company Ltd, makers of product Y has prepared its budget for 2018, based on two activity levels of 80% and 100% with production units of 2800 and 3500 units respectively.

The budget is as follows

                                               80%                                         100%

Sales                                    $224000                                   $2800000

Direct material                       84000                                    105000

Direct labour                         50400                                     63000

Production overhead              48800                                    53000

The company’s actual result for the period is as follows

Sales in Units                                          3150 units

Sales                                                        $252000

Direct material                                        $ 94000

Direct Labour                                          $ 60000

Overhead (60% of amounts is fixed)      $ 50000

Required

Prepare a flexible budget for a production level of 70% and 90% and compare the flexible budget at 90% level of activity with the actual result. (25)

Question 2

A business supplies the following figures about its activities

Fixed                                                     $ 300000

Variable cost                                         $ 20 per unit

Forecast output (sales)                           20000 units

Selling price                                           $50 per unit

Required

Illustrate by means of breakeven chart

  1. The breakeven point
  2. The profit at full capacity
  3. The margin of safety
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Answer(1) :

We are given with the following data for 100% activity levels

Sales Direct Material Direct Labour Production Overhead $280,000 $105,000 $63000 $53000

Let us prepare a budget plan for 70% and 90% activity levels from this table

170% 90% Actual figures Calculations Remarks Category Units 2450 =(3500*70/100) 3150 = (3500*90/100) Sales Direct Material Di

Some of the workings:

Sales/Unit =   $2800000/3500 = $80

Direct material costs =  105000/3500 =$30

Fixed costs = $32,000

Variable costs = $21,000/3500 = $6

Answer(2) :

Breakeven point calcualtion is given by =(Fixed Cost/Contribution margin per unit)

=> Contribution margin = Selling price/Unit - Variable Cost per unit

=> Contibution margin = $50 - $20 =$30

=> Breakeven point = $300000/$30 = 10000 units

Profit at full capacity is given by:

Sales at full capacity for 20,000 units = $50*20,000 =$1,000,000

Less (Variable cost for 20,000 units) = 20*$20,000 = $40,000

Less (Fixed Costs) =$30,000

=> Profit at full capacity =$1,000,000 - $40,000 - $30,000 = $30,000

Margin of safety calcualtion is given by Actual sales - Breakven sales

=> Margin of safety = $1,000,000 - $50*10000 = $50,000